OpenStudy (anonymous):

A substance has a half-life of 60,000 years and an initial mass of 1,000 grams. How many years will pass before only 250 grams of the substance's parent material is left? 30,000 90,000 120,000 240,000

OpenStudy (anonymous):

I say A

OpenStudy (anonymous):

how so

OpenStudy (anonymous):

half of 60,000

OpenStudy (anonymous):

I have 2 more questions, can you help?

OpenStudy (anonymous):

Sure :)

OpenStudy (anonymous):

You are given a box of 100 silver dollars, all facing "heads" up. You are instructed to shake the box 3 times; after each shake, you will remove all the dollars that are "heads up" before shaking again. You may keep all the dollars that are still tails up following the third shake. How many dollars will you most likely get to keep? 6 12 25 50

OpenStudy (anonymous):

probability of a dollar remaining a heads = 1/2 Expected no. of tails remaining therefore after 1st shake = 1/2 * 100 = 50 similarity after second shake 25

OpenStudy (anonymous):

I say 25 idk

OpenStudy (anonymous):

Yep

OpenStudy (anonymous):

A rubidium isotope has a half-life of almost 50 billion years. Given that scientists estimate Earth's age to be 4.6 billion years, what is the most likely percentage of parent to daughter isotopes of this element currently existing on Earth? Less than 10 percent 25 percent 50 percent More than 75 percent

OpenStudy (anonymous):

http://openstudy.com/study#/updates/51db482ae4b0d9a104d98861 Here this guy explains the good

OpenStudy (anonymous):

:o

OpenStudy (anonymous):

thank u

OpenStudy (anonymous):

Any time :)